Determine which of the ordered pairs
(0, 1), (2, 0), (0, –1), (–8, 5)
are solutions for the equation x + 2y = 2.

Answer by mananth(16946)

plug values given .
if they are equal then it is a solution

(0, 1), 0+2=2 , 2=2
(2, 0), 2+0=2, 2=2
(0, –1), 0+(-2)=2, 2 not equal to 2 Not a solution
(–8, 5) -8+10=2, 2=2

equation x + 2y = 2.